A mass m = 0.1 gram is hanged on a string inside a bus. The bus is accelerating at a constant rate of 5 m/s^2. At what angle will the mass deflect from the vertical? (in degrees)?


1
Expert's answer
2022-03-29T12:29:23-0400

Answer

Angle is "\theta"

So

"Masin\theta=Mg"

So angle


"\theta=sin^{-1}(g/a) \\=sin^{-1}(9.8/5)"